OpenStudy (anonymous):

HELP: The formula A=P(1+r)^2 gives the amt A in dollars, that P dollars will grow to in 2 years at interest rated r (where r is given as a decimal), using the compound interest? What interest rate will cause 6000 to grow into 6678.15 in two years? I am stuck. A=P(1+r)^2 6678.15=6000(1+r)^2 6678.15/6000=(1+r)^2 1.11302=(1+r)^2 I am stuck at this point!

OpenStudy (anonymous):

(1 + r)^2 = 1.11302 1 + r = sqrt(1.11302) r = sqrt(1.11203) - 1 its now calculator time

OpenStudy (anonymous):

sorry thats sqrt(1.11302) - 1 = .055 or 5.5 %

OpenStudy (anonymous):

A=P(1+r)^2 -> (1 + r )^2 = A/P -> 1 + r = sqrt (A/P) => r = sqrt [(A/P) - 1 ] Thus rate = sqrt (6678.15 / 6000 ) - 1 = 11.3%

OpenStudy (anonymous):

i think your calculation is wrong chlorophyll

OpenStudy (anonymous):

Yes, I'm testing why. Thanks

OpenStudy (anonymous):

Sorry! sqrt (6678.15 / 6000 ) = 1.055 r = 1.055 -1 = .055 = 5.5%

OpenStudy (anonymous):

thats ok - good job

OpenStudy (anonymous):

I really learn a lot from other solvers like you :)

OpenStudy (anonymous):

glad to hear it - thats what this site is for